In line with the overall development of the health infrastructure, the state governments across the country provides a grant of Rs.6,00,000 each year to various laboratories across various organisations to conduct cancer research, where the costs of operating the research lab is Rs.3,00,000. Prima Donna, which is a leading alcohol beverage manufacturing company, offers funding of Rs. 8,00,000 each year to do beer research, but researching on beer would increase costs by Rs. 50,000.

1. What is the opportunity cost of conducting cancer research? Break this into implicit and explicit costs.

2. What option will the department choose? What are the opportunity costs of this choice?


1
Expert's answer
2022-02-11T08:32:32-0500

1. What is the opportunity cost of conducting cancer research? Break this into implicit and explicit costs.

Opportunity Cost = sum of Explicit Cost (Cancer Research) + Implicit Cost (forgone profits from beer research) to Pursue.

Explicit cost is Rs. 300,000 while implicit cost is Rs. 450,000.

The Explicit cost of the cancer research is give 300000 while the implicit cost of the forgone profits from Prima Donna is given as 450000

Therefore, the Opportunity Cost = 300000 + 450000 = 750000

Thus, the opportunity cost of conducting cancer research is Rs. 750,000


The profits are calculated as picture below:



2. What option will the department choose? What are the opportunity costs of this choice?

If the department's selection is based on profit, the logical option is Beer Research. The economic profit if they choose bear research is Rs. 150000.

Opportunity cost

The opportunty cost is 650000 of conducting beer research. Given by:

Opportunity Cost = sum of Explicit Cost (beer Research) + Implicit Cost (forgone profits from cancer research) to Pursue.

Opportunity Cost = 350000 + 300000 = Rs. 650000